OpenStudy (yusmin):

red marbles 9 blue marbles and 6 green marbles. what is the probability of choosing a marble that is not green?

OpenStudy (yusmin):

can someone please help me?

OpenStudy (zzr0ck3r):

1-minus the prob that you will get a green

OpenStudy (calculusxy):

How many red marbles are there? Is there any missing info?

OpenStudy (yusmin):

7 marbles sorry

OpenStudy (calculusxy):

What's the total number of marbles?

OpenStudy (yusmin):

22

OpenStudy (calculusxy):

Subtract the number of green marbles from 22 because you don't want to choose green marbles.

OpenStudy (yusmin):

So it will be 16

OpenStudy (calculusxy):

Yes. That 16 will be the numerator and the 22 (total) will be the denominator. You would have something like \( \bf \large \frac{16}{2}\)

OpenStudy (calculusxy):

Sorry I meant 16/22

OpenStudy (calculusxy):

\(\large \bf \frac{16}{22}\) You could simplify this more by dividing by 2 to get \(\large \frac{8}{11}\). However, I am not sure whether they want it as a fraction or as a decimal or percent. Do you have the answer choices?

OpenStudy (yusmin):

I don't but it says it's right :)

OpenStudy (calculusxy):

Ok that's great!

OpenStudy (yusmin):

Thank you so much

OpenStudy (calculusxy):

You're welcome :)

OpenStudy (zzr0ck3r):

Note that another way of doing this is to see that there is a \(\frac{6}{22}\) chance you get a green one, so there is a \(1-\dfrac{6}{22}=\frac{22}{22}-\frac{6}{22}=\frac{16}{22}=\frac{8}{11}\) Sometimes this method is easier.